Osborn Manufacturing uses a predetermined overhead rate of $18.80 per direct labor-hour. This predetermined rate was based on a cost formula that estimates $236,880 of total manufacturing overhead for an estimated activity level of 12,600 direct labor-hours. The company actually incurred $233,000 of manufacturing overhead and 12,100 direct labor-hours during the period. Required: 1. Determine the amount of underapplied or overapplied manufacturing overhead for the period. 2. Assume that the company's underapplied or overapplied overhead is closed to Cost of Goods Sold. Would the journal entry to dispose of the underapplied or overapplied overhead increase or decrease the company’s gross margin? By how much?
Answer #1

1) Applied overhead = 12100*18.80 = 227480

Actual overhead = 233000

Under applied overhead = 233000-227480 = 5520

2a) Journal entry

No General Journal Debit Credit
Cost of goods sold 5520
Manufacturing overhead 5520

By this entry, Cost of goods sold increase and gross margin decrease by $5520